x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6 x/3-x/2=6

 x/3-x/2=6
 x/3-x/2=6
 x/3-x/2=6
 x/3-x/2=6
 
 x/3-x/2=6
 x/3-x/2=6
 x/3-x/2=6

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

6x – 9 -2(1+x) = x -9

6x -9 -2 – 2x = x – 9

6x – 2x – x = -9 + 9 + 2

3x = 2

x = 2/3

Thanks